# A Principle of Targeted Intervention for Multi-Agent Reinforcement Learning

### v3

@@ 6 Conclusion
Summary. We incorporate multi-agent influence diagrams (MAIDs) as a principled framework into designing and analyzing the targeted intervention paradigm in MARL. Our Pre-Strategy Intervention (PSI) approach as an implementation of the targeted intervention paradigm, applied to a single targeted agent, maximizes the causal effect on achieving a composite desired outcome that integrates both the primary task goal and an additional desired outcome. Furthermore, the relevance graphs of MAIDs offer theoretical insights into the solvability of MARL learning paradigms under various interaction paradigms. Experiments verify the effectiveness of the PSI.
@@
    Designing the optimal criteria for selecting the number of targeted agents and appointing proper agents with various types to complete different tasks in various scenarios. The information-theoretic measure of empowerment jaques2019social ; klyubin2005empowerment offers a potential quantitative basis for such criteria.
@@
    Integrating advanced reasoning modules (e.g. large language models and the broader generative AI) to enhance the capability of the PSI on the targeted agent in online learning and online adaptation to unknown agent teammates and unseen situations (wang2024open, ).
@@
    The principle proposed in this paper can be used to design new MARL interaction paradigms.
@@ Appendix A Notation
Table 1: Summary of notations used in this paper.
@@
$U_{tot}^{task},U_{tot}^{sec}$ Utility of the primary task goal and the secondary goal, respectively.
@@ B.1 Multi-agent Team with Human Feedback
Our work connects to the growing field of leveraging human guidance within multi-agent systems [45]. A prominent related area is reinforcement learning from human feedback (RLHF), where human input, such as preferences or demonstrations, guides agent policy optimization [46, 47], often proving essential for tasks with complex or subjective goals for which explicit reward design is difficult [47, 48, 49]. RLHF is a specific instance within the broader scope of human-instructed multi-agent systems, which utilize diverse inputs like direct commands [50], demonstrations [51, 52], policy shaping [48], or language instructions [53] to steer agent behaviors. Incorporating such human expertise can significantly enhance system performance [46, 54, 55, 56], interpretability [11, 53, 57, 58], and alignment [47, 59]. These methods typically treat human feedback as a data source to directly train an agent’s policy, often in a model-free manner.
@@ B.2 Goal-Conditioned MARL
Our work relates to goal-conditioned reinforcement learning (GCRL), where agents learn policies conditioned on achieving diverse goals, such as reaching specific states [60, 61], matching images [62], or following language instructions [63]. While concepts similar to subgoals like roles [27], skills [64], or subtasks [65] are used in MARL for task decomposition or coordination. The foundational approach in this area is the option framework in hierarchical reinforcement learning (HRL) [66], where “options” (temporally extended sub-policies) are used to achieve subgoals.
@@ B.3 Intrinsic Reward Method
A common approach to realize guidance in MARL is through maximizing intrinsic rewards during learning. These internally generated signals can encourage diverse behaviours motivated by curiosity [67, 68], novelty [69, 70], social influence [7], empowerment [9], learned reward functions [8, 71] or utilizing prior knowledge from large language models [23, 53].
While our Pre-Strategy Intervention is implemented via an intrinsic reward, our core distinction lies in its purpose and scope. Unlike methods such as those in  [8, 23], which applied intrinsic rewards to all agents to encourage emergent coordination, our Pre-Strategy Intervention strategically directs the signal only to a single targeted agent. Furthermore, standard intrinsic rewards are generally designed to solve local challenges like sparse rewards via using heuristic metrics like novelty. In contrast, our guidance signal is explicitly designed to solve a system-level problem: reaching a desired outcome for the entire multi-agent system.
@@ B.4 Environment and Mechanism Design
Our work relates to concepts from environment and mechanism design. Environment design typically involves structuring or modifying environmental configurations to guide agent behaviors towards a targeted outcome [72, 73, 74]. In contrast, our approach does not directly reconfigure the static environment. Rather, it focuses on intervening on a targeted agent. Consequently, the behavior of the intervened agent changes, altering the effective environment dynamics experienced by other agents. This allows us to shape emergent team behavior by modifying agent interaction paradigms rather than fixed environmental rules.
Our work can also be viewed through the lens of mechanism design, which focuses on designing the “rules of games” such that a desired outcome is attained [75, 76]. Our proposed targeted intervention paradigm can be categorized into this context as a mechanism that is applied to specific agents designed to steer the MARL learning process towards equilibria that satisfy both the primary task goal and a specified additional desired outcome.
@@ B.5 Probabilistic Graphical Models for Multi-Agent Games
Graphical models based on Bayesian networks were introduced to represent dependencies, relevance, and relationships among variables in multi-agent games by Koller and Milch [17]. This formulation has since inspired advancements in both algorithmic innovations for solving games [77, 78] and empirical analysis [79]. More recently, Hammond et al. [80] incorporated causal graphical models into multi-agent reasoning, extending multi-agent influence diagrams (MAIDs) to the causal domain. Their work introduced the concept of pre-policy intervention, allowing for queries such as “What if some agents commit to a policy before others make their decisions?” This framework has become a foundational tool in subsequent research on theory of mind [81, 82], reinforcement learning [83], and causal modeling of agents [42]. To maintain consistency with the MAID framework, we rename “pre-policy intervention” to “pre-strategy intervention” in this paper.
Our work builds directly on the causal framework proposed by Hammond et al. [80]. However, whereas their work introduced pre-policy intervention as a general tool for causal reasoning in game theory, our work is the first to adapt and apply this concept to solve a core challenge in MARL: achieving coordination by guiding a multi-agent system towards a desired outcome. We leverage MAIDs not just for analysis, but as a practical design tool to construct an MARL interaction paradigm. We design Pre-Strategy Intervention as the core mechanism within the targeted intervention paradigm, using it to guide the MARL learning process towards a desired, high-performing outcome.
@@ C.1 An Example of MAIDs
We exemplify MAIDs through a two-agent scenario from [17].
@@ C.2 Additional Definitions
Definition C.1 (Active Path [84]).
Let $\mathcal{G}$ be a Bayesian Network (BN) structure, and let $X_{1}-X_{2}-\dots-X_{n}$ represent an undirected path in $\mathcal{G}$. Let $E$ be a subset of nodes in $\mathcal{G}$ (the evidence set). The path $X_{1}-\dots-X_{n}$ is active given evidence $E$ if the following conditions hold:
@@
    No other nodes along the path is in $E$.
@@ D.1 Pre-Strategy Intervention
(b) To prove that there exists a pre-strategy maximizing the causal effect, we observe that the second term on the right-hand side of Equation (2) is constant. Therefore, maximizing the first term is equivalent to maximizing the causal effect.
To analyze how an agent’s strategy influences the final outcome, we can decompose the probability of achieving the desired outcome $U_{tot}=u^{*}$. Using the law of total probability and the Markov property of the Bayesian network [17], we can marginalize out the decision variable $D$ and its parents $Pa(D)$. This decomposition explicitly isolates the agent’s decision rule, $P_{\mathcal{M}[\sigma]}(d\mid\mathbf{pa}_{D})$, which is the component that our intervention targets:
@@ Appendix F Multi-Agent Influence Diagrams and Relevance Graphs in Implementation
$D_{h}$$U_{h}$$D_{\text{pre}}$$Z_{h}$$D_{a}$$U_{a}$
$D_{h}$$D_{a}$
(a) Simultaneous Moves.
[…]
$D_{h}$$D_{a}$
(b) Sequential Moves.
Figure 7: MAIDs and relevance graphs for MARL interaction paradigms. The figures illustrate the structure for (a) Simultaneous moves and (b) Sequential moves. The variable “ZZ” here indicates an intrinsic reward to implement an additional desired outcome.
$D^{t}_{h}$$D^{t}_{a}$$D^{t}_{\text{pre}}$$Z^{t}$$S^{t}$$U_{h}^{t}$$U^{t}$$U_{a}^{t}$$S^{t+1}$
@@
Figure 8: MAIDs for pre-strategy intervention approach implemented in experiments, in sequential-move sequential decision makings, their corresponding Markov games and the relevance graphs for the Markov games. The variable $Z^{t}$” here indicates an intrinsic reward to implement an additional desired outcome.
In this section, we provide a more detailed illustration of the multi-agent influence diagram (MAID) structure that represents the practical implementation of our proposed Pre-Strategy Intervention. This shows how the intervention, designed to guide a single targeted agent towards an additional desired outcome, is integrated into that agent’s decision-making process within the MAID framework. We present these implementation-specific MAIDs and their corresponding relevance graphs for both one-shot and sequential decision-making scenarios.
F.1 Pre-Strategy Intervention Implementation in One-Shot Decision Making
Figure 7 illustrates the MAID structures representing our implemented Pre-Strategy Intervention for one-shot decision-making in both simultaneous and sequential moves. These diagrams illustrate how the pre-decision node ($D_{pre}$) and the guidance signal/additional desired outcome variable ($Z$) (represented as an intrinsic reward) directly inform the targeted agent’s decision ($D_{h}$). Furthermore, an edge from a variable representing the additional desired outcome $Z$ to agent’s utility node $U_{h}$ indicates that the agent’s utility now explicitly incorporates this additional desired outcome.
F.2 Pre-Strategy Intervention Implementation in Sequential Decision Making
Figure 8(a) illustrates the MAID structures and its implementation for applying Pre-Strategy Intervention for sequential decision making with simultaneous moves. Similarly, Figure 8(b) shows these for sequential-move scenarios. These diagrams show the temporal extension, where the intervention influences decisions at each relevant timestep.
Appendix G Representing Markov Games through Multi-Agent Influence Diagrams
This appendix details how a standard team reward Markov game can be equivalently represented as a Multi-Agent Influence Diagram (MAID).
Markov game [26] is a popular mathematical model to describe the multi-agent sequential decision process across various real-world applications [85, 86, 87]. For succinct description, we only consider the team reward Markov game with a finite episode length, as described in Definition G.1.
Definition G.1 (Markov Game [36]).
@@ Appendix G Representing Markov Games through Multi-Agent Influence Diagrams
A team reward Markov game can be rendered as a MAID, because we can match variables between these two models. In more details, both models’ agent sets are $\mathcal{I}$; $\mathcal{S}$ is associated with chance variables $\mathcal{X}$; $\mathcal{A}$ is associated with decision variables $\mathcal{D}$; $T$ is associated with conditional probability distributions $Pr$; $\pi$ is associated with decision rules $\delta$; and $\mathbb{E}_{\pi,Pr}[\sum_{t=0}^{L}R(s_{t},a_{t})]$ is associated with the expected utility shown in Equation 1. The objective of a team reward Markov game $\max_{\pi}\mathbb{E}_{\pi,Pr}[\sum_{t=0}^{L}R(s_{t},a_{t})]$ is equivalent to reaching a Nash equilibrium (see Definition 2.2), given that agents share total utility variables as team rewards.
@@ Appendix H Implementation Details
The implementation is mainly based on JaxMARL [88]. Our experiments were run on NVIDIA RTX 4090 and A100 GPUs. A typical training run for the Hanabi environment takes approximately one hour on these systems.
This section details the architecture of our proposed method and the baselines used for comparison. Subsequently, we describe how the “additional desired outcome” is defined and implemented within the MPE and Hanabi environments respectively.
@@ H.1 Implementation of Our Method
This entire architecture, including the agent’s policy, the GNN module and the pre-policy module, is trained jointly end-to-end to maximize a composite shaping reward. The composite shaping reward is the sum of the extrinsic task reward and the intrinsic reward derived from the additional desired outcome. As established in Section 4.3, this training process is theoretically sound, as maximizing the intervention’s causal effect requires observing the total team utility ($U_{tot}$), which directly aligns with the use of a shared team reward in many previous MARL works.
@@
Targeted Agent Selection. For the experiments presented in this paper, we manually select a targeted agent in each environment. This experimental design allows for a clear and controlled analysis of the mechanism’s impact and mirrors practical scenarios where a human operator might manually choose a specific agent to intervene on.

### v4

@@ 6 Conclusion
Summary. We incorporate multi-agent influence diagrams (MAIDs) as a principled framework into designing and analyzing the targeted intervention paradigm in MARL. Our Pre-Strategy Intervention (PSI) approach as an implementation of the targeted intervention paradigm, applied to a single targeted agent, maximizes the causal effect on reaching a composite desired outcome that integrates both the primary task goal and an additional desired outcome. Furthermore, the relevance graphs of MAIDs offer theoretical insights into the solvability of MARL learning paradigms under various MARL interaction paradigms. Experiments verify the effectiveness of the PSI.
@@
    Designing the optimal criteria for selecting the number of targeted agents and appointing proper agents with various types to complete different tasks. The information-theoretic measure of empowerment jaques2019social; klyubin2005empowerment offers a potential quantitative basis for such criteria.
@@
    Integrating advanced reasoning modules (e.g. large language models and the broader generative AI) to enhance the capability of the PSI on the targeted agent in online learning and online adaptation to unknown agent teammates and unseen situations (wang2024open).
@@
    Designing MARL algorithms with asynchronous updates and mixed learning paradigms inspired by the targeted intervention paradigm (Footnote 4).
@@ Appendix A Notation
Table 1: Summary of key notations used in this paper.
@@
$U_{task},U_{sec}$ Utility of the primary task goal and the secondary goal, respectively.
@@ B.1 Multi-agent Team with Human Feedback
Our work connects to the growing field of leveraging human guidance within multi-agent systems [dorri2018multi]. A prominent related area is reinforcement learning from human feedback (RLHF), where human input, such as preferences or demonstrations, guides agent policy optimization [griffith2013policy, ouyang2022training], often proving essential for tasks with complex or subjective goals for which explicit reward design is difficult [ouyang2022training, hadfield2016cooperative, akrour2014programming]. RLHF is a specific instance within the broader scope of human-instructed multi-agent systems, which utilize diverse inputs like direct commands [wu2022survey], demonstrations [carroll2019utility, hester2018deep], policy shaping [hadfield2016cooperative], or language instructions [hu2023language] to steer agent behaviours. Incorporating such human expertise can significantly enhance system performance [griffith2013policy, zhang2020kogun, liu2024integrating, chen2025hierarchical], interpretability [hu2022human, hu2023language, cui2023adversarial, yan2024efficient], and alignment [ouyang2022training, zhang2024multi]. These methods typically treat human feedback as a data source to directly train an agent’s policy, often in a model-free manner.
@@ B.2 Goal-Conditioned MARL
Our work relates to goal-conditioned reinforcement learning (GCRL), where agents learn policies conditioned on achieving diverse goals, such as reaching specific states [andrychowicz2017hindsight, liu2022goal], matching images [nair2018visual], or following language instructions [luketina2019survey]. While concepts similar to subgoals like roles [wang2020roma], skills [yang2019hierarchical], or subtasks [yang2022ldsa] are used in MARL for task decomposition or coordination. The foundational approach in this area is the option framework in hierarchical reinforcement learning (HRL) [sutton1999between], where “options” (temporally extended sub-policies) are used to achieve subgoals.
@@ B.3 Intrinsic Reward Method
A common approach to realize guidance in MARL is through maximizing intrinsic rewards during learning. These internally generated signals can encourage diverse behaviours motivated by curiosity [pathak2017curiositydrivenexplorationselfsupervisedprediction, zheng2021episodic], novelty [kayal2025impact, tao2020novelty], social influence [jaques2019social], empowerment [mohamed2015variational], learned reward functions [du2019liir, zheng2018learning] or utilizing prior knowledge from large language models [wang2025m3hf, hu2023language].
While our Pre-Strategy Intervention is implemented via an intrinsic reward, our core distinction lies in its purpose and scope. Unlike methods such as those in  [du2019liir, wang2025m3hf], which applied intrinsic rewards to all agents to encourage emergent coordination, our Pre-Strategy Intervention strategically directs the signal only to a single targeted agent. Furthermore, standard intrinsic rewards are generally designed to solve local challenges like sparse rewards via using heuristic metrics like novelty. In contrast, our guidance signal is explicitly designed to solve a system-level problem: reaching a desired outcome for the entire multi-agent system.
@@ B.4 Environment and Mechanism Design
Our work relates to concepts from environment and mechanism design. Environment design typically involves structuring or modifying environmental configurations to guide agent behaviours towards a targeted outcome [zhang2009general, reda2020learning, gao2023constrained]. In contrast, our approach does not directly reconfigure the static environment. Rather, it focuses on intervening on a targeted agent. Consequently, the behaviour of the intervened agent changes, altering the effective environment dynamics experienced by other agents. This allows us to shape emergent team behaviour by modifying agent interaction paradigms rather than fixed environmental rules.
Our work can also be viewed through the lens of mechanism design, which focuses on designing the “rules of games” such that a desired outcome is attained [nisan1999algorithmic, cai2013understanding]. Our proposed targeted intervention paradigm can be categorized into this context as a mechanism that is applied to specific agents designed to steer the MARL learning process towards equilibria that satisfy both the primary task goal and a specified additional desired outcome.
@@ B.5 Probabilistic Graphical Models for Multi-Agent Games
Graphical models based on Bayesian networks were introduced to represent dependencies, relevance, and relationships among variables in multi-agent games by Koller and Milch [koller2003multi]. This formulation has since inspired advancements in both algorithmic innovations for solving games [vickrey2002multi, kearns2013graphical] and empirical analysis [wellman2006methods]. More recently, Hammond et al. [hammond2023reasoning] incorporated causal graphical models into multi-agent reasoning, extending multi-agent influence diagrams (MAIDs) to the causal domain. Their work introduced the concept of pre-policy intervention, allowing for queries such as “What if some agents commit to a policy before others make their decisions?” This framework has become a foundational tool in subsequent research on theory of mind [ward2023honesty, ward2024reasons], reinforcement learning [macdermott2024measuring], and causal modeling of agents [kenton2023discovering]. To maintain consistency with the MAID framework, we rename “pre-policy intervention” to “pre-strategy intervention” in this paper.
Our work builds directly on the causal framework proposed by Hammond et al. [hammond2023reasoning]. However, whereas their work introduced pre-policy intervention as a general tool for causal reasoning in game theory, our work is the first to adapt and apply this concept to solve a core challenge in MARL: achieving coordination by guiding a multi-agent system towards a desired outcome. We leverage MAIDs not just for analysis, but as a practical design tool to construct an MARL interaction paradigm. We design Pre-Strategy Intervention as the core mechanism within the targeted intervention paradigm, using it to guide the MARL learning process towards a desired, high-performing outcome.
@@ C.1 An Example of MAIDs
We exemplify MAIDs through a two-agent scenario from [koller2003multi].
@@ C.2 Additional Definitions
Definition C.1 (Active Path [pearl2014probabilistic]).
Let $\mathcal{G}$ be a Bayesian Network (BN) structure, and let $X_{1}-X_{2}-\dots-X_{n}$ represent an undirected path in $\mathcal{G}$. Let $E$ be a subset of variables in $\mathcal{G}$ (the evidence set). The path $X_{1}-\dots-X_{n}$ is active given evidence $E$ if the following conditions hold:
@@
    No other variables along the path is in $E$.
@@ D.1 Pre-Strategy Intervention
(b) To prove that there exists a pre-strategy maximizing the causal effect, we observe that the second term on the right-hand side of Eq. (2) is constant. Therefore, maximizing the first term is equivalent to maximizing the causal effect.
To analyze how an agent’s strategy influences the final outcome, we can decompose the probability of achieving the desired outcome $U_{tot}=u^{*}$. Using the law of total probability and the Markov property of the Bayesian network [koller2003multi], we can marginalize out the decision variable $D$ and its parents $Pa(D)$. This decomposition explicitly isolates the agent’s decision rule, $P_{\mathcal{M}[\sigma]}(d\mid\mathbf{pa}_{D})$, which is the component that our intervention targets:
@@ Appendix F Multi-Agent Influence Diagrams and Relevance Graphs in Implementation
$D^{t}_{h}$$D^{t}_{a}$$D^{t}_{\text{pre}}$$Z^{t}$$S^{t}$$U_{h}^{t}$$U_{team}^{t}$$U_{a}^{t}$$S^{t+1}$
@@
Figure 7: MAIDs for pre-strategy intervention approach implemented in experiments, in sequential-move sequential decision makings, their corresponding Markov games and the relevance graphs for the Markov games. The figures illustrate the structure for (a) Simultaneous moves and (b) Sequential moves. The variable $Z^{t}$” here indicates an intrinsic reward to implement an additional desired outcome, attributed to the the single targeted agent’s individual utility variable $U_{h}^{t}$. This means the reachability of the additional desired outcome can be observed by the single targeted agent during learning.
In this section, we provide a more detailed illustration of the multi-agent influence diagram (MAID) structure that represents the practical implementation of our proposed Pre-Strategy Intervention. This shows how the intervention, designed to guide a single targeted agent towards an additional desired outcome, is integrated into that agent’s decision-making process within the MAID framework. We present these implementation-specific MAIDs and their corresponding relevance graphs for sequential decision-making scenarios.
Figure 7(a) illustrates the MAID structures and its implementation for applying Pre-Strategy Intervention for sequential decision making with simultaneous moves. Similarly, Figure 7(b) shows these for sequential-move scenarios. These diagrams illustrate how the pre-decision variable ($D^{t}_{pre}$) and an additional desired outcome variable ($Z^{t}$) (represented as an intrinsic reward) directly inform the targeted agent’s decision ($D^{t}_{h}$) at each relevant timestep. Furthermore, an edge from a variable representing the additional desired outcome $Z^{t}$ to the targeted agent’s individual utility variable $U^{t}_{h}$ indicates that it can observe the reachability of the additional desired outcome during learning.
Appendix G Representing Team Reward Markov Games as Multi-Agent Influence Diagrams
This appendix details how a team reward Markov game can be equivalently represented as a Multi-Agent Influence Diagram (MAID).
Markov games [littman1994markov] is a popular mathematical model to describe the multi-agent sequential decision process across various real-world applications [qiu2021multi, wang2021multi, zhang2024improving]. For succinct description, we only consider the team reward Markov game with a finite episode length, as described in Definition G.1.
Definition G.1 (Team Reward Markov Game [littman2001value]).
@@ Appendix G Representing Team Reward Markov Games as Multi-Agent Influence Diagrams
A team reward Markov game can be represented as a MAID, because we can match variables between these two models. In more details, both models’ agent sets are $\mathcal{I}$; $\mathcal{S}$ is associated with chance variables $\mathcal{X}$; $\mathcal{A}$ is associated with decision variables $\mathcal{D}$; $T$ is associated with conditional probability distributions $Pr$; $\pi$ is associated with decision rules $\delta$; and $\mathbb{E}_{\pi,T}[\sum_{t=0}^{L}R(s_{t},a_{t})]$ is associated with the expected utility shown in Eq. (1), with representing $R(s_{t},a_{t})$ as the team utility variable $U_{team}^{t}$.
For the self-organization, each agent’s individual utility variable in the MAID can represent the equal contribution of the team utility variable $U_{team}^{t}$ that implies a team reward. According to [wang2020shapley], this is equivalent to the case where each agent’s individual utility variable directly represents the shared team reward $R(s_{t},a_{t})$. Therefore, maximizing the objective of a team reward Markov game $\max_{\pi}\mathbb{E}_{\pi,T}[\sum_{t=0}^{L}R(s_{t},a_{t})]$ is equivalent to reaching a Nash equilibrium (NE) (Definition 2.2). For the targeted intervention, the team utility variable is defined as the total utility variable: $U_{team}^{t}:=U_{tot}^{t}=U_{task}^{t}+U_{sec}^{t}$, where $U_{task}^{t}$ is shared among agents and $U_{sec}^{t}$ is only attributed to the targeted agent. Thus, the targeted agent’s individual utility variable is defined as $U_{task}^{t}+U_{sec}^{t}$, while other agents’ are defined as $U_{task}^{t}$. Therefore, maximizing each agent’s individual utility leads to the preferred NE among multiple equilibria of the primary task utility. For the global intervention with an additional outcome, both $U_{task}^{t}$ and $U_{sec}^{t}$ are shared among agents. According to [wang2020shapley], each agent’s individual utility variable is equally defined as $U_{task}^{t}+U_{sec}^{t}$. For the global intervention without an additional outcome, the situation is the same as the self-organization.
@@ Appendix H Implementation Details
The implementation is mainly based on JaxMARL [flair2023jaxmarl]. Our experiments were run on NVIDIA RTX 4090 and A100 GPUs. A run of experiments for the Hanabi environment typically completes in about one hour on these systems, while a run of experiments for the MPE environment only requires 5 minutes. The following paragraphs detail the architecture of our proposed method and the baselines used for comparison. Subsequently, we describe how the “additional desired outcome” is defined and implemented within the MPE and Hanabi environments respectively.
@@ H.1 Implementation of Our Method
This entire architecture, including the agent’s policy, the GNN module and the pre-policy module, is trained jointly end-to-end to maximize a composite shaping reward. The composite shaping reward is the sum of the extrinsic task reward and the intrinsic reward derived from the additional desired outcome. As established in Section 4.3, this training process is theoretically sound, as maximizing the intervention’s causal effect requires observing the total utility variable ($U_{tot}$), which directly aligns with the use of a shared team reward in many previous MARL works.
@@
Targeted Agent Selection. For the experiments presented in this paper, we manually select a single targeted agent in each environment. This experimental design allows for a clear and controlled analysis of the mechanism’s impact and mirrors practical scenarios where a human operator might manually choose a specific agent to intervene on.